# Quantum cohomology of projective bundles

### v3

@@ 3.2. Informal proof
arising from the Stirling approximation does not converge in an adic topology, because the exponent contains [^6] This is also the case with the original operator $\Delta_{(W,e_{\lambda}^{-1})}$, but this was not a problem because the coefficient of $z^{-1}$ in $\log(\Delta_{(W,e_{\lambda}^{-1})})$ is nilpotent. both positive and negative powers of $z$. This converges in the standard topology to power series in $z$ and $z^{-1}$ (infinite in both directions) with coefficients in analytic functions of $\lambda$ near $\lambda=e$, but we are not sure whether this fact would be useful. The only properties of $\Delta_{W}^{\lambda}$ we use in the rigorous argument below (e.g. (3.4)) can be interpreted in terms of the well-defined logarithm $\log(\Delta^{\lambda}_{W})\in z^{-1}H^{*}(B)\otimes\mathbb{C}[\lambda,\lambda^{-1},\log\lambda][\![z]\!]$ and the convergence issue does not exist.
3.3. Rigorous proof
We explain how to make the above argument rigorous, using the $\lambda^{-1}$-adic topology. We follow the strategy explained in Remark 2.6. Let $\tilde{e}_{\lambda}(\cdot)=\sum_{i\geq 0}\lambda^{-i}c_{i}(\cdot)$ be the modified equivariant Euler class and let $\widetilde{\Delta}_{W}^{\lambda}:=\Delta_{(W,\tilde{e}_{\lambda}^{-1})}$ denote the QRR operator (2.19) of the $(W,\tilde{e}_{\lambda}^{-1})$-twist. We have $\widetilde{\Delta}_{W}^{\lambda}=1+O(\lambda^{-1})$. We also set $\widetilde{J}_{V}^{\lambda}(\tau):=J_{V}^{\lambda}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$; this does not contain positive powers of $\lambda$ and gives the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory. We have
$\Delta_{W}^{\lambda}$ $=e^{(\operatorname{rank}(W)\theta(\lambda)+c_{1}(W)\log\lambda)/z}\widetilde{\Delta}_{W}^{\lambda}\quad\text{with $\theta(\lambda):=\lambda\log\lambda-\lambda+\frac{\log\lambda}{2}z$}$
$J_{V}^{\lambda}(\tau)$ $=e^{(\log\lambda)(c_{1}(V)-z\partial_{c_{1}(V)})/z}\widetilde{J}_{V}^{\lambda}(\tau)$
where $\partial_{c_{1}(V)}$ means the directional derivative along $c_{1}(V)$ in the $\tau$-space. The second equation follows from the Divisor Equation. Combining these equations, we have
(3.8) $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)\quad\text{with $f(\lambda):=\operatorname{rank}(\mathcal{Q})\theta(\lambda)-(\log\lambda)z\partial_{c_{1}(V)}$.}$
Since $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{\Delta}_{V}^{\lambda}=1$ (by the exact sequence (3.2)), $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)$ is a $\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5. This is a miniversal slice, i.e. $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)|_{Q=\lambda^{-1}=0}=z+\tau+O(z^{-1})$. Therefore, we can apply Lemma 2.8 to this slice, and obtain a differential operator $\widetilde{F}(\lambda)\in\sum_{i=0}^{s}\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]z\partial_{\tau^{i}}$ such that $\widetilde{F}(\lambda)|_{Q=\lambda^{-1}=0}=0$ and
Mimicking the previous argument in §3.2, we consider applying the differential operator $e^{f(\lambda+z\partial_{t})/z}e^{\widetilde{F}(\lambda+z\partial_{t})/z}$ to $J_{B\times\mathbb{P}^{N-1}}(\tau,t)$; the first factor $e^{f(\lambda+z\partial_{t})/z}$ was added in view of the difference (3.8) between $\Delta^{\lambda}_{\mathcal{Q}}J_{V}^{\lambda}$ and $\widetilde{\Delta}^{\lambda}_{\mathcal{Q}}\widetilde{J}_{V}^{\lambda}$. To work with $\lambda^{-1}$-power series, we subtract from $f(\lambda+z\partial_{t})$ the first few terms of its Taylor expansion in $z\partial_{t}$ and define:
$f_{\rm reg}(\lambda,z\partial_{t}):=f(\lambda+z\partial_{t})-f(\lambda)-\operatorname{rank}(\mathcal{Q})(\log\lambda)z\partial_{t}.$
This belongs to $\mathbb{C}[z\partial_{t},z\partial_{c_{1}(V)}][\![\lambda^{-1}]\!]$ and $f_{\rm reg}(\lambda,z\partial_{t})|_{\lambda^{-1}=0}=0$. Therefore, Lemma 2.7 and the QRR Theorem 2.5 imply that
is a $\mathbb{C}[\![Q,\tau,t,\lambda^{-1}]\!]$-valued point of the $(\mathcal{Q}(1),\tilde{e}_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$. By calculations similar to, but more cumbersome than, the previous section §3.2, we arrive at the following result (cf. (3.6)).
$\widetilde{I}^{\lambda}(\tau,t)=\sum_{k\geq 0}e^{pt/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}\left(1+\frac{p+\epsilon+cz}{\lambda}\right)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$
where $\prod_{\epsilon}$ is the product over the Chern roots $\epsilon$ of $\mathcal{Q}$. For this calculation, we used (3.9) and the following identities:
$\widetilde{\Delta}^{\lambda}_{\mathcal{Q}(1)}$ $=e^{(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+p)-\theta(\lambda)-p\log\lambda)+c_{1}(\mathcal{Q})\log(1+\frac{p}{\lambda}))/z}\widetilde{\Delta}^{\lambda+p}_{\mathcal{Q}},$
$\widetilde{\Delta}_{\mathcal{Q}}^{\lambda+kz}/\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}$ $=e^{-\left(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+kz)-\theta(\lambda))+c_{1}(\mathcal{Q})\log(1+\frac{kz}{\lambda})\right)/z}\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+\epsilon+cz),$
$\ J_{V}^{\lambda+x}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ $=e^{-c_{1}(\mathcal{Q})\log(1+\frac{x}{\lambda})/z}e^{-\log(1+\frac{x}{\lambda})\partial_{c_{1}(V)}}\widetilde{J}_{V}^{\lambda+x}(\tau).$
The first identity follows from $\Delta^{\lambda}_{\mathcal{Q}(1)}=\Delta^{\lambda+p}_{\mathcal{Q}}$, the second from (3.4), and the third from the Divisor Equation. Finally we make the change of variables $Q^{d}\to Q^{d}\lambda^{c_{1}(\mathcal{Q})\cdot d}=Q^{d}\lambda^{-c_{1}(V)\cdot d}$, $q^{k}\to q^{k}\lambda^{\operatorname{rank}(\mathcal{Q})k}$ for $\widetilde{I}^{\lambda}$ and obtain a point $I^{\lambda}$ of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone. This gives the same result as (3.6). The non-equivariant limit $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ exists and gives the $I$-function $I_{\mathbb{P}(V)}$ lying on the Givental cone of $\mathbb{P}(V)$. The proof of Theorem 1.1 is now complete.
@@ 4.1. Quantum $D$-modules
We fix notation and definitions concering the quantum $D$-modules of $\mathbb{P}(V)$ and $V$.
@@ 4.1.1. Quantum $D$-module of $\mathbb{P}(V)$
Let $\hat{\tau}\in H^{*}(\mathbb{P}(V))$ denote the parameter of the quantum cohomology of $\mathbb{P}(V)$. Following Notation 3.1, we write $(q,Q)$ for the Novikov variables of $\mathbb{P}(V)$, which are dual to the decomposition $H^{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}p\oplus H^{2}(B,\mathbb{Z})$. The quantum $D$-module of $\mathbb{P}(V)$ is the module
@@
induced by the Poincaré pairing of $\mathbb{P}(V)$. We note that $\nabla_{q\partial_{q}}=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}})$. We also note that we defined $\operatorname{QDM}(\mathbb{P}(V))$ over $\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$ instead of over $\mathbb{C}[z][\![q,Q,\hat{\tau}]\!]$: it is possible because of the homogeneity in Gromov-Witten theory and the fact that both $z$ and $q$ have positive degrees.
@@ 4.1.2. Quantum $D$-module of $V$
Let $\tau\in H^{*}_{S^{1}}(V)$ denote the parameter of the equivariant quantum cohomology of $V$. Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(B)$ as before; here we assume that $\phi_{0}=1$. Then $\{\phi_{i}\lambda^{k}\}_{0\leq i\leq s,k\geq 0}$ gives a $\mathbb{C}$-basis of $H^{*}_{S^{1}}(V)$. We write $\tau=\sum_{i=0}^{s}\sum_{k=0}^{\infty}\tau^{i,k}\phi_{i}\lambda^{k}$ with $\tau^{i,k}$ being $\mathbb{C}$-valued coordinates. The equivariant quantum $D$-module of $V$ is the module
@@ 4.2. Fourier transformation for quantum $D$-modules
Recall the Fourier transformation (1.2) on the level of the Givental spaces from the introduction.
@@
where $\mathcal{H}_{V,\rm pol}$, $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ are the polynomial Givental spaces given in (2.17), (2.15). (Note that $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ is defined over the Novikov ring $\mathbb{C}[\![q,Q]\!]$.)
@@
Let $V\to B$ be as in Theorem 1.1. There exist a mirror map $\hat{\tau}=\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[q][\![Q,\boldsymbol{\tau}]\!]$ and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules which satisfy the following properties:
@@
    $\hat{\tau}(\tau)$ is homogeneous of degree $2$ and $\operatorname{FT}$ is homogeneous of degree 00 with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(\mathbb{P}(V))$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2c_{1}(TV)\cdot d=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
@@
    ${\operatorname{QDM}_{S^{1}}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{FT}\quad}$$\scriptstyle{M_{V}(\tau)}$${\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{\mathbb{P}(V)}(\hat{\tau}(\tau))}$${\mathcal{H}_{V,\rm pol}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\phantom{aa}}}$${\mathcal{H}_{\mathbb{P}(V),\rm pol},}$
where we set $\nabla^{\prime}=\hat{\tau}^{*}\nabla$.
Remark 4.3.
The pull-back $\nabla^{\prime}=\hat{\tau}^{*}\nabla$ is defined in the standard way as follows:
$\nabla^{\prime}_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{\mathbb{P}(V)}\star_{\hat{\tau}(\tau)})+\mu_{\mathbb{P}(V)}.$
@@
and these vectors with $0\leq i\leq s$, $10\leq k\leq r-$ 1 form a basis of $H^{*}(\mathbb{P}(V))$. The mirror map $\hat{\tau}(\tau)$ is then determined by $M_{\mathbb{P}(V)}(\hat{\tau}(\tau))1=1+z^{-1}\hat{\tau}(\tau)+O(z^{-2})$. The formula (4.2) with $i=k=0$ implies that $\hat{\tau}(\tau)|_{Q=q=0}=\kappa(\tau)$ and $\Upsilon(\tau)|_{Q=q=0}=1$. Similarly, a direct calculation shows that $(M_{V}(\tau)1)^{^}\,|_{Q=0,\tau^{*,r}=\tau^{*,r+1}=\cdots=0}\equiv 1+z^{-1}\kappa(\tau)+O(z^{-2})$ modulo $(\tau^{*,\leq r-1})^{2}$. Part (7) follows from these computations.
$M_{V}(\tau)\phi_{i}\lambda^{k}$ is homogeneous of degree $\deg(\phi_{i}\lambda^{k})$. It is easy to see that its Fourier transform $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,$ is also homogeneous of the same degree. This implies the homogeneity of the Birkhoff factors $M_{\mathbb{P}}(V)(\hat{\tau}(\tau))$, $\Upsilon(\tau)$, $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ and hence of the mirror map $\hat{\tau}(\tau)$. The homogeneity implies that $\hat{\tau}(\tau)$ and $\Upsilon(\tau)$ depend polynomially on $q$ (for each fixed power of $Q$ and $\tau$), i.e. $\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!]$, $\Upsilon(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$.
@@
which shows that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $10\leq k\leq r-$ 1 and, since the multiplication by $\lambda$ corresponds to $z\nabla^{\prime}_{q\partial_{q}}$ under $\operatorname{FT}$, that
@@
Hence for $k=rn+m$ with $10\leq m\leq r-$ 1, we have
@@
    $\pi_{*}(\kappa(\lambda^{i}))=0$ for $20\leq i\leq r-$ 2 and $\pi_{*}(\kappa(\lambda^{r-1}))=1$.

### v4

@@ 3.1. Outline
We note that the modified QRR operator (3.4) arising from the Stirling approximation does not converge in an adic topology, because the exponent contains [^8] This also holds for the original operator $\Delta_{(W,e_{\lambda}^{-1})}$; however this is not an issue because the coefficient of $z^{-1}$ in $\log(\Delta_{(W,e_{\lambda}^{-1})})$ is nilpotent. both positive and negative powers of $z$. This converges in the standard topology to power series in $z$ and $z^{-1}$ (infinite in both directions) with coefficients in analytic functions of $\lambda$ near $\lambda=e$, but we are not sure whether this fact would be useful. The properties of $\Delta_{W}^{\lambda}$ we used in the proof (e.g. (3.5)) can be interpreted in terms of the well-defined logarithm $\log(\Delta^{\lambda}_{W})\in z^{-1}H^{*}(B)\otimes\mathbb{C}[\lambda,\lambda^{-1},\log\lambda][\![z]\!]$ and the convergence issue does not exist.
@@ 4.1. Quantum $D$-modules
We fix notation and definitions concerning the quantum $D$-modules of $\mathbb{P}(V)$ and $V$.
@@ 4.1.1. Quantum $D$-module of $\mathbb{P}(V)$
Let $\hat{\tau}\in H^{*}(\mathbb{P}(V))$ denote the bulk-deformation parameter of the quantum cohomology of $\mathbb{P}(V)$. Following Notation 3.1, we write $(q,Q)$ for the Novikov variables of $\mathbb{P}(V)$, which are dual to the decomposition $H^{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}p\oplus H^{2}(B,\mathbb{Z})$. The quantum $D$-module of $\mathbb{P}(V)$ is the module
@@
[…]
@@ 4.1.2. Quantum $D$-module of $V$
Let $\tau\in H^{*}_{S^{1}}(V)$ denote the bulk-deformation parameter of the equivariant quantum cohomology of $V$. Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(B)$ as before; here we assume that $\phi_{0}=1$. Then $\{\phi_{i}\lambda^{k}\}_{0\leq i\leq s,k\geq 0}$ gives a $\mathbb{C}$-basis of $H^{*}_{S^{1}}(V)$. We write $\tau=\sum_{i=0}^{s}\sum_{k=0}^{\infty}\tau^{i,k}\phi_{i}\lambda^{k}$, where the $\tau^{i,k}$ are $\mathbb{C}$-valued coordinates, and we use $\boldsymbol{\tau}$ to denote the infinite set $\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$ of coordinates. The equivariant quantum $D$-module of $V$ is the module
@@ 4.2. Fourier transformation for quantum $D$-modules
Recall the Fourier transformation (1.3) on the level of the Givental spaces from the introduction:
@@
where $\mathcal{H}_{V,\rm pol}$, $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ are the polynomial Givental spaces given in (2.17), (2.15). (We define $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ over the completed Novikov ring $\mathbb{C}[\![q,Q]\!]$.)
@@
Let $V\to B$ be as in Theorem 1.1. There exist a mirror map $\hat{\tau}=\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[q][\![Q,\boldsymbol{\tau}]\!]$ and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules, such that by defining the pull-back connection $\nabla^{\prime}=\hat{\tau}^{*}\nabla$ via the standard formulae
$\nabla^{\prime}_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{\mathbb{P}(V)}\star_{\hat{\tau}(\tau)})+\mu_{\mathbb{P}(V)},$
the following properties hold:
@@
    $\hat{\tau}(\tau)$ is homogeneous of degree $2$ and $\operatorname{FT}$ is homogeneous of degree 0 with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(\mathbb{P}(V))$ and the following degrees of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2c_{1}(TV)\cdot d=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
@@
    ${\operatorname{QDM}_{S^{1}}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{FT}\quad}$$\scriptstyle{M_{V}(\tau)}$${\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{\mathbb{P}(V)}(\hat{\tau}(\tau))}$${\mathcal{H}_{V,\rm pol}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}^\scriptstyle{\widehat{\phantom{aa}}}$${\mathcal{H}_{\mathbb{P}(V),\rm pol}.}$
@@
and these vectors with $0\leq i\leq s$, $0\leq k\leq r-1$ form a basis of $H^{*}(\mathbb{P}(V))$. The mirror map $\hat{\tau}(\tau)$ is then determined by $M_{\mathbb{P}(V)}(\hat{\tau}(\tau))1=1+z^{-1}\hat{\tau}(\tau)+O(z^{-2})$. The formula (4.2) with $i=k=0$ implies that $\hat{\tau}(\tau)|_{Q=q=0}=\kappa(\tau)$ and $\Upsilon(\tau)|_{Q=q=0}=1$. Similarly, a direct calculation shows that $(M_{V}(\tau)1)^{^}\,|_{Q=0,\tau^{*,r}=\tau^{*,r+1}=\cdots=0}\equiv 1+z^{-1}\kappa(\tau)+O(z^{-2})$ modulo $(\tau^{*,\leq r-1})^{2}$. Part (7) follows from these computations.
$M_{V}(\tau)\phi_{i}\lambda^{k}$ is homogeneous of degree $\deg(\phi_{i}\lambda^{k})$. It is easy to see that its Fourier transform $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,$ is also homogeneous of the same degree. This implies the homogeneity of the Birkhoff factors $M_{\mathbb{P}}(V)(\hat{\tau}(\tau))$, $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ and hence of $\Upsilon(\tau)=\nabla^{\prime}_{\tau^{0,0}}\Upsilon(\tau)$ and the mirror map $\hat{\tau}(\tau)$. The homogeneity implies that $\hat{\tau}(\tau)$ and $\Upsilon(\tau)$ depend polynomially on $q$ (for each fixed power of $Q$ and $\tau$), i.e. $\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!]$, $\Upsilon(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$.
@@
which shows that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $0\leq k\leq r-1$ and, since the multiplication by $\lambda$ corresponds to $z\nabla^{\prime}_{q\partial_{q}}$ under $\operatorname{FT}$, that
@@
Hence for $k=rn+m$ with $0\leq m\leq r-1$, we have
@@
    $\pi_{*}(\kappa(\lambda^{i}))=0$ for $0\leq i\leq r-2$ and $\pi_{*}(\kappa(\lambda^{r-1}))=1$.